The vapor pressure of the liquid HF is measured at different temperatures. The following vapor pressure data are obtained: Temperature 270.6K and 287.5K, Pressure 324.5 mmHg and 626.9 mmHG. Calculate the enthapy of vaporization ( delta H vap ) in kJ/mol for this liquid.
Answer #1

1) Let us use the Clausius-Clapeyron Equation:

ln (P1 / P2) = (?H / R) (1/T2 - 1/T1)

with the following values:

P1 = 324.5 mmHg T1 = 270.6 K
P2 = 626.9 mmHg   T2 = 287.5 K

2) Set up equation with values:

ln (324.5/ 626.9) = (x / 8.31447) (1/287.5 minus 1/270.6)

-0.658 = (x / 8.31447) (-0.000217)

5.4708 = 0.0001217x (notice I got rid of the negative signs)

x =44953 J/mol = 44.95 kJ/mol